OpenStudy (wgibson08):

Which expression can be used to find the difference of the polynomials? (10m – 6) – (7m – 4)

OpenStudy (krazonix):

I like that you stuck a mathematics question in the [English] category. Kudos to you :) (10m - 6) - (7m - 4) = 0 Now solve the brackets: 10m - 6 - 7m + 4 = 0 3m - 2 = 0 Finally, solve for x! 3m = 2 m = 2/3